### v1

# Retina Vision Transformer (RetinaViT): Introducing Scaled Patches into Vision Transformers

Abstract
Humans see low and high spatial frequency components at the same time, and combine the information from both to form a visual scene. Drawing on this neuroscientific inspiration, we propose an altered Vision Transformer architecture where patches from scaled down versions of the input image are added to the input of the first Transformer Encoder layer. We name this model Retina Vision Transformer (RetinaViT) due to its inspiration from the human visual system. Our experiments show that when trained on the ImageNet-1K dataset with a moderate configuration, RetinaViT achieves a 3.3% performance improvement over the original ViT. We hypothesize that this improvement can be attributed to the inclusion of low spatial frequency components in the input, which improves the ability to capture structural features, and to select and forward important features to deeper layers. RetinaViT thereby opens doors to further investigations into vertical pathways and attention patterns.
Keywords:
Biologically-inspired neural networks Multi-scale visual processing Vision Transformers
1 Introduction
Work on neural networks in machine learning has often drawn inspiration from brain science, from the model of McCulloch and Pitts [18] that led to the Perceptron [21], to CNNs [15], attention [24], and even consciousness [1]. However one insight from neuroscience that has been less explored is that of processing order in the human visual cortex [20].
1.1 Inspiration
The inspiration of this paper is a difference between the state-of-the-art backbone computer vision models and human vision. All backbone computer vision models use one single image, at one single scale, as the input to the model, while human vision processes the visual scene at different resolutions at the same time. Humans have two photoreceptor systems within the retina, namely the rod and cone systems, for low and high resolution vision, respectively [20]. On a more macro level, human vision processes low and high spatial frequency components in the scene via two different neural pathways, at two different speeds [23]. In other words, humans perceive both low and high spatial frequency signals and combine the information therein in visual processing.
The literature has not discussed, to the best of our knowledge, why only one single scale is used extensively in the input of deep neural network-based vision models. This might be due to implementation costs, and the intuition that information in lower resolution versions of an image is already entailed in the higher resolution versions.
The authors, on the contrary, hypothesize that including information from different resolutions of the same image may improve model performance. We propose an architecture which uses a pyramid of scaled images, as illustrated in fig. 1, instead of just the one image at one scale as the input, to test this assumption.
(a) “Image pyramid” constructed from the original resolution image (base of the pyramid). Patches from images at each level are used as input to RetinaViT.
(b) Size of a patch $(16\times 16)$ relative to the size of the “base-level”, i.e., original resolution $(224\times 224)$ image, giving a total of 196 patches.
(c) Size of a patch $(16\times 16)$ relative to the size of the “mid-level” scaled $(64\times 64)$ image, giving a total of 16 patches.
(d) Size of a patch $(16\times 16)$ comprises the entire “top-level” scaled $(16\times 16)$ image, i.e., it is just the whole image shrunk to a single patch.
Figure 1: Example of an image pyramid with edge lengths of 224, 128, 64, 32, and 16 pixels. For illustrative purposes the subfigures are not to scale. Original image from [8].
Notice that, unlike some previous work [6], this image pyramid has no semantic meaning in itself, but is simply a stack of the same image at a range of different resolutions.
Also, drawing on another biological analogy, the authors refer to the area covered by a patch as its receptive field [10]. The receptive field of a neuron is the area where sensory signals can elicit an action potential from the neuron, for example the dendritic fields of different types of ganglion cells in the retina [14].
In the context of this paper, we use it to refer to the area of the input image covered by a patch, because any change in this area would cause the embedding of the patch to change. Specifically, as illustrated in fig. 1, patches from a smaller scale image cover a larger area of the scene, and thus have a larger receptive field.
1.2 Feasibility of Introducing Multiple Scales into Model Input
In the field of computer vision, since the resurgence of deep learning, there have been two main backbone models, namely the Convolutional Neural Network (CNN) [15] and the Vision Transformer (ViT) [9].
CNN performs multiple convolution operations on overlapping patches of the input image, and uses max pooling to forward the presence of the most important features from one layer to the next. Notice the spatial structure of the input image is preserved in CNNs, i.e., the information at a certain location of the input image on one layer will be forwarded to a less accurate, but similar location in the next layer.
ViT, on the other hand, performs one single convolution on patches of the input image to obtain the embeddings of the patches, and relies on the self-attention mechanism to determine the effective order of the patches to attend to. In other words, the spatial structure of the input image is not preserved per se, only encoded into positional embeddings and added to the embeddings of the patches as a single vector.
It is more difficult to introduce scaled versions of the original image into CNNs, because from the second layer onwards, convolutions are performed on the feature maps output from the previous layers, instead of scaled versions of the original image. It is not easy to reconcile both into the input of the deeper layers.
ViT, on the other hand, allows this introduction much more easily. By flattening the patches from the input image into a one dimensional vector, ViT decouples the patches from their spatial arrangement in the input image. The patches are no longer constrained by their spatial arrangement, but become in some sense a “bag of patches”.
By the same token, we can remove the constraint of the hierarchical structure of the image pyramid, and concatenate patches from all its levels into one single vector, and use this vector as the input to the Transformer Encoder. Notice spatial information is not lost, it is still preserved in the positional embeddings added to the patch embeddings. Only the structural constraint is removed.
Based on the above, this paper proposes an architecture where patches from an image pyramid containing the input image in a range of resolutions are concatenated into one single vector, and used as the input to the ViT.
Due to its biological inspiration, this model is named Retina Vision Transformer (Retina ViT).
In the remainder of this paper, we demonstrate empirically in  section 4 that this modification improves the performance of the model on a standard benchmark dataset. In  section 5 we discuss some of the theoretical implications of this modification.
2 Related Work
2.1 Smaller stride and patch size
An investigation into the convolution kernel used in the patch extraction process in ViT [29] finds that a smaller patch size and stride improves the top-1 accuracy of ViT by 1-2%. RetinaViT does not reduce the patch size, but does use a stride of half the patch size in patch extraction.
2.2 Training ViT at different scales
CrossViT [7] combines two Transformers using two different patch sizes with cross attention. This setup shares the idea of using patches that cover receptive fields of different sizes, but the two Transformers are not sharing weights, and thus are not capturing scale invariant features as much. CrossFormer [27] attempts to capture scale invariant features by extracting embeddings from patches of four different sizes, centered at the same location. Notice this setup also has the potential of reflecting characteristics of human vision.
More recently, ResFormer [25] trains ViT with images scaled to different resolutions, and uses global-local positional embeddings to maintain cross resolution consistency. Each image is scaled to three resolutions, but these are then fed into the model independently, instead of as a concatenated patch sequence. FlexiViT [2] randomises patch size to allow one single model to process a wide range of patch sizes. Notice that randomising image resolution and patch size are both randomising the relative receptive field size of each patch, as compared to the size of the scene.
2.3 Introducing CNN traits into ViT
ViT has fewer inductive biases than CNN. Many models have explored the effect of reintroducing some of the inductive biases of CNN into ViT.
2.3.1 Spatial Dimensionality Reduction
Pooling-based Vision Transformer (PiT) [11] shows that reducing the spatial dimensionality as the network goes deeper is beneficial to ViT as well. Pyramid Vision Transformer (PVT) [26] uses a progressively shrinking pyramid with inter-layer spatial-reduction attention, to perform dense prediction tasks.
2.3.2 Convolutions in ViT
Convolutional Vision Transformers (CvT) [28] introduce shift, scale, and distortion invariance into ViT by: a) partitioning ViT into a hierarchy of three stages, and reducing the number of tokens in the convolutional embedding layer at the beginning of each stage, to achieve downsampling, and b) replacing the linear projection prior to the attention layer in a normal Transformer block with convolutional projections. Notice that the standard embedding layer in a ViT is already convolutional, so a) is more about staging and applying the embedding layer more often.
2.3.3 Local Attention Windows
Swin Transformer [16], among many other models, limits self attention to non-overlapping local windows, and merges such windows as the network goes deeper. This is also similar to the idea of max pooling downsampling in CNNs.
3 Retina Vision Transformer
Figure 2: The architecture of RetinaViT. This figure is adapted from the illustration of the architecture of the original ViT [9], and an image from the ImageNet-1k dataset [8].
3.1 Model Architecture
The architecture of RetinaViT, as shown in fig. 2, is very close to that of the original ViT, except for the following changes.
3.1.1 Adding Scaled Patches
RetinaViT adds patches from an image pyramid, with different resolutions of the original image, to the input of the Transformer Encoder. The original image with a resolution of $(224\times 224)$ is scaled down into four lower resolutions $(128\times 128)$, $(64\times 64)$, $(32\times 32)$, and $(16\times 16)$, then the whole hierarchy of five is flattened into one sequence of patches, and patch embeddings are extracted with the same convolution kernel therefrom. Notice the first scale down is from $(224\times 224)$ to $(128\times 128)$, instead of $(112\times 112)$, because 112 is not divisible by patch size 16, and the authors do not wish to introduce further deviations from the patch extraction process of the original ViT.
3.1.2 Scaled Average Positional Embedding
The positional embedding of RetinaViT is extended from the 2-D sincos (sincos2d) embedding implementation in the big_vision library [4]. In the original ViT architecture, with an input size of $(224\times 224)$ and a patch size of $(16\times 16)$, the positional embeddings form a $(14\times 14)$ matrix. RetinaViT first backfills the embedding vectors into a $(224\times 224)$ matrix, where each embedding vector from the $(14\times 14)$ matrix covers a $(16\times 16)$ area. Then for each scaled patch, RetinaViT calculates the weighted average embedding vector of all pixels covered by the particular patch in the original image, and scales its norm to the square root of the relative edge length of the receptive field of the scaled patch compared to that of the original patch.
For example, in fig. 3, a patch from the $(64\times 64)$ resolution image covers a $(56\times 56)$ area in the original $(224\times 224)$ image, and its positional embedding is calculated as the average positional embedding of the pixels in that $(56\times 56)$ area, scaled to a norm of $\sqrt{56/16}$.
Figure 3: A patch from the $(64\times 64)$ image, denoted with the green box, covers a larger area of the scene than the patches from the original image, denoted with white boxes.
Just as positional embeddings are used to preserve the information of the location of the patches in ViT, the norms of the average positional embeddings in RetinaViT are used to preserve the information of the relative receptive field size. As illustrated in fig. 4, the scaled average positional embedding of a patch points to a location in a 3-dimensional image pyramid, instead of a location on a 2-dimensional image.
(a) The positional embedding of a patch from the original image represents a 2-D vector.
(b) The positional embedding of a patch from the image pyramid represents a 3-D vector.
Figure 4: Comparing the positional embeddings of an image and that of an image pyramid.
3.1.3 Smaller Stride
RetinaViT uses the same patch size of $(16\times 16)$ as the original ViT, but uses a stride of half the patch size (8), instead of the whole of it. With only the first two modifications, the model does not exhibit much improvement over the original ViT, but when stride is halved, the model exhibits the final result in this paper, i.e., a 3.3% increase in top-1 prediction accuracy, over the original ViT.
The authors believe that there is a synergy between the first two modifications, and a smaller stride. As mentioned in section 2.1, using a smaller patch size and stride alone improves the performance of ViT by 1-2%, so the rest of the performance improvement can be attributed to the first two modifications.
3.2 Configurations
Due to budget constraints, the authors followed the setup in [3], and did not perform extensive hyperparameter tuning. The authors expect that with more tuning, the performance of RetinaViT could be further improved.
The results in this paper are from a RetinaViT model based on the ViT-S/16 variant, trained for 90 epochs on ImageNet-1K, using the same hyperparameters as those defined in the vit_s16_i1k configuration of the big_vision library [4]. Specifically, the ViT-S/16 variant uses a hidden state size of 384, network depth of 12 layers, MLP layers of size 1536, and 6 attention heads. The training batch size is 1024. The same augmentation as those used in [3] are also adopted, namely random horizontal flips, RandAugment, and Mixup.
The implementation of RetinaViT can be found at https://github.com/yuyangshu/RetinaViT.
4 Experimental Evaluation
Model Validation set
val real v2
Original 76.5% 83.1% 64.4%
RetinaViT 79.8% 85.7% 68.5%
Difference +3.3% +2.6% +4.1%
Table 1: Top-1 prediction accuracies of RetinaViT, compared to those of the original ViT.
table 1 contains the experiment results of RetinaViT, as compared to that from the original ViT with the same setup. Across the three test datasets, RetinaViT achieves an average improvement of 3.3%. Notice the accuracies for the original ViT are those quoted in [3]. In our experiments when we ran the code for the original ViT, the accuracy figures we obtained were slightly lower, but we use the published results in case our replication missed some details of the experiments in the original paper.
fig. 5 shows how the prediction accuracies of RetinaViT and ViT changed, as training proceeded. Notably, the stability of the prediction accuracy of RetinaViT has not changed much, compared to that of the original ViT.
Figure 5: Top-1 prediction accuracy graph of RetinaViT (blue), on the val dataset, compared to the original ViT implementation (orange).
In terms of added costs, RetinaViT only adds more patches to the initial layer of the Transformer Encoder, and thus only causes a 5.3% increase in the total number of parameters. More specifically, since the images are usually scaled down by a factor of 2, the theoretical limit of the total number of patches fed into the first layer of RetinaViT over that of ViT is the geometric series with a coefficient of 1, and a common ratio of $1/4$, equivalent to a 33.3% increase. In our implementation, the first scale down is from 224 to 128, thus the actual increase in the number of parameters in the first layer is 43.4%.
Notice that with the same configuration, the time cost of training RetinaViT increased from 6.5 hours on a v3-8 TPU node, to 30 hours on a v3-32 TPU node. However, the authors believe that this can be mitigated because
  1. 1.
    The positional embedding calculations is highly reusable, as the embedding process is independent of the content of the images.
  2. 2.
    The scaling of the input image should be performed in the processing pipeline, instead of repeatedly inside the model.
The authors have not implemented these optimisations for a cleaner comparison with the original ViT implementation.
5 Discussion
The architectural changes in RetinaViT have a few theoretical implications.
5.1 Conceptual Expansion of Input Dimensionality
With the attention mechanism, Transformer effectively transforms an input sentence into a parse tree [5]. In recurrent models such as Long Short-term Memory (LSTM) [13], tokens in the input are fed into the model sequentially by their order of appearance in the sentence. But with attention, it is possible for a token with a high attention score to be effectively processed before preceding tokens. This is equivalent to allowing Transformers to move freely on a parse tree of the input sentence, and process tokens in an arbitrary order.
Similarly, by introducing scaled patches into ViT, RetinaViT transforms the input image, which is effectively the bottom layer of an image pyramid, into the full image pyramid, and allows attention to decide the order in which to process the patches. This is illustrated in fig. 6.
(a) An example parse tree for text input with a potential attention path.
(b) A flattened image pyramid with a potential attention path.
Figure 6: Potential attention paths in Transformer on text and RetinaViT.
This effectively expands the dimensionality of the input from two to three, as illustrated in fig. 4. Notice this expansion is conceptual only, since the format of the input has not changed, and there is no new information added to the input vector. The input to the Transformer Encoder is still a sequence of embedded patches, and all information in the patches is derived from the original image.
5.2 More Capable of Capturing Features
Compared to the original ViT, RetinaViT is also more capable of capturing features, in the following ways:
5.2.1 Scale-invariant Features
Traditionally, computer vision models, such as Scale-invariant Feature Transform (SIFT) [17], look for scale-invariant features that are preserved when the scale of the image is changed. This allows the recognition of the same features, and subsequently objects, in a close up as well as a further away image.
CNN is not innately scale-invariant [30]. Although it may be able to capture scale-invariant features to some extent, this is more due to the fact that each convolution layer has multiple kernels, and some of them may capture the same scale-invariant features at different scales. The original ViT is not scale-invariant either, since the patches are all from an image of one single scale.
RetinaViT is closer to CNN than ViT, because its input contains the same image at multiple different scales. This means that each scale-invariant feature is guaranteed to be seen by RetinaViT $n$ times, where $n$ is the number of layers in the input image pyramid. As a result, RetinaViT is stronger than both CNN and ViT in recognizing scale-invariant features, and has a higher chance of detecting the same objects at various scales.
It also has a greater potential to capture cross-layer relationships, because given that all patches are embedded with the same convolution kernels, patches from different scales but containing the same scale-invariant features, at the same location of the scene, should have closer patch embeddings and positional embeddings.
5.2.2 Semantic Unit in Vision
Another factor is that RetinaViT has a higher chance of capturing the basic semantic units in vision. Language models are token based, which means their input has already been preprocessed into semantically meaningful units. On the other hand, it is more difficult to define a semantically meaningful unit for vision models, and subsequently preprocess them into model inputs. Intuitively, blobs of a similar colour or texture, or areas enclosed by a contour, could constitute basic units in vision. But irrespective of the precise definition, the size and shape of such basic units could vary considerably. By including patches of the same image at different scales, RetinaViT has a higher chance of capturing a semantically meaningful basic unit in its input. Using a stride of half the patch size also helps on this front, because some features could happen to lie on the border of two patches extracted with equal patch size and stride.
5.2.3 Low and High Spatial Frequency Components
Introducing the ability to process both low and high spatial frequency components of the visual scene into the model is the main motivation of RetinaViT. Patches at a smaller scale carry more information from the low spatial frequency components of the input image, whereas patches at a larger scale carry more information from the high spatial frequency components. Therefore, RetinaViT is capable of capturing features that are more prominent in the low spatial frequency components of the image, with the addition of the smaller scale patches.
5.3 On Inductive biases
The authors of the original paper on ViT state that fewer inductive biases are present in the ViT setup, and explain the improved performance of ViT over CNN with “large scale training trumps inductive bias” [9].
On one hand, we agree that fewer inductive biases are helpful in some ways. In the case of RetinaViT, it is exactly the absence of a strict structure on the input, which is an inductive bias, that allows the introduction of patches from smaller scale images. But on the other hand, we believe that inductive biases are not necessarily bad, and it is possible to reintroduce some useful inductive biases back into models with inherently few inductive biases, such as ViT, to improve performance. In other words, requiring large scale training is a shift in paradigm, but that does not prohibit models under the new paradigm from using inductive biases where appropriate.
Notice that RetinaViT has not hand-crafted any attention patterns, since how attention scores are assigned to each patch is still learned during the training process. Therefore no extra inductive bias is introduced on this front.
One inductive bias that is indeed reintroduced into RetinaViT, in a very subtle way, is that as the network goes deeper, only the more important features are forwarded to the next layer.
In CNN, as the network goes deeper, the size of the input to each layer shrinks, whereas the receptive field size of each convolution kernel increases. In the final layer, only the most important features are preserved after the Max Pooling processes throughout network.
In some other models, such as the spiking neural network framework Assembly Calculus [19], there is a Random Projection and Cap primitive, where one group of neurons randomly connect to downstream neurons, and only the connections that are the most consistent with Hebbian plasticity are preserved. This is also a selection process, with a cap on information capacity.
In RetinaViT, patches from multiple scales are added to the input of the Transformer Encoder, so there is more information in the first layer than the original ViT. But since the size of the Transformer blocks has not changed, there is effectively a capping process, where the most important information from patches from all five scales are selected, and forwarded to deeper layers.
6 Future Work
The authors have formulated several hypotheses based on RetinaViT, and plan to investigate these in future work.
6.1 Emergence of Vertical Pathways
Learning part-whole relationships, or how features extracted from fine-grained details such as texture can be bound with those from higher level characteristics such as contour, has always been a focus of studies. For example, CapsNet [22] uses agreement between activity vectors of capsules, and GLOM [12] uses spatially local columns, to bind features from different scales to a single processing pathway.
In CNN, due to the preservation of the spatial arrangement of features across layers, binding of low and high level features is achieved by design. In ViT, because of the flexibility in processing order provided by the attention mechanism, there is not such a binding mechanism.
In RetinaViT, there is no explicit binding mechanism either, but the authors hypothesize that including patches from different scales can enable the emergence of different vertical pathways inside RetinaViT. It is possible for RetinaViT to form a pathway where the initial information comes from patches containing low spatial frequency components, and only interact with another pathway that initiates from patches containing high spatial frequency components deeper down the network. The authors plan to look for such pathways in future studies.
6.2 Attention Patterns
CNNs have a less-discussed inductive bias, that processing starts from more local details, and ends at the higher level features and characteristics. This is contrary to how humans process visual input. As discussed in [23], humans process low spatial frequency components quicker than high spatial frequency components. This helps perceive the “gist” of the scene as a whole first, before turning to more nuanced matters.
With RetinaViT, since low and high spatial frequency components are fed into the Transformer Encoder at the same time, it is possible for low spatial frequency components to be attended to, before high spatial frequency components are. The authors plan to investigate such patterns in future work as well.
7 Conclusion
In this paper, we investigated the hypothesis that providing information from the same image in multiple scales to a neural network may allow it to learn a better model for a visual task. To test this idea in a controlled way, we devised an approach whereby we supply all patches from a hierarchy of scaled images, referred to as an image pyramid, as input to the standard Vision Transformer model. We showed experimentally that this approach can improve performance on a benchmark dataset. Several directions for future work arise from the discussion of our approach, which we plan to investigate further.